# Demonstrating Wheeled Lab: Modern Sim2Real for Low-cost, Open-source Wheeled Robotics

### v1

@@ V-B Elevation
Background. On unstructured and uneven terrain, an elevation map is essential for identifying what areas in the scene are traversable. However, traversability is strongly dependent on agent morphology and dynamic state. Along with the ground geometry, the platform’s ground clearance, center-of-mass height, wheel size, maximum torque, suspension, and momentum, all affect how to safely traverse uneven terrain. Poor traversal can result in catastrophic failures such as rollovers and motor (or engine) stalling.
@@
Task. To show that the trained elevation-based policy is capable of geometric reasoning, we construct a scene with two types of elevation features which contribute to the overall elevation map: walls and ramps. Crucially, the incorporation of ramps sets the task apart from sole obstacle avoidance, since ramps are features with height but are still traversable with the correct approach strategy. The starting and goal positions are placed on opposite sides of the testing area with the elevation features placed in between. The platform must safely traverse the scene to arrive at the goal position, avoiding obstacles and ascending ramps when locally feasible.
@@
Setup. States provided to the policy are goal-relative position, orientation, and a local ($2.5\times 2.5$ meters) body-centric elevation map. A motion capture system is used for creating the global map from which we sample local maps. While the platform is capable of creating onboard maps [45], a configurable indoor environment requires higher mapping resolution than available for method evaluation.
Results.
Figure 6: Three-dimensional view of global elevation map. $\boldsymbol{\pi}_{\text{elev}}$ (yellow/blue) and $\overline{\boldsymbol{\pi}}_{\text{elev}}$ (purple/white) trajectories are visualized over it with elevation values denoted using color. Note that the policies only have access to their local $2.5\times 2.5$ m maps during evaluation.
We find that the baseline primarily deviates from any elevation features to approach the goal. When the baseline does ascend the ramp, it quickly falls off the sides.
@@ Results.
However, $\boldsymbol{\pi}_{\text{elev}}$ is capable of both traversing the ramp safely and navigating through the subsequent obstacles to reach the goal. Similar to $\overline{\boldsymbol{\pi}}_{\text{drift}}$, the lower number of simulated agents gives the baseline policy significantly fewer opportunities to explore correct trajectories in a narrow range of safe states.
V-C Visual
@@
Figure 7: Captured trajectories of visual policy in real experiments. The trajectories are obtained from the motion capture system. Videos of each setting and trial are included in the supplementary materials.
@@
TABLE III: Visual Policy Results. This demonstrates the effectiveness of image augmentations for Sim2Real generalization. While we expect CNN to be more capable of learning, MLP presents better generalization capability. Videos of each run are included in the supplementary materials.
Background. Visual navigation traditionally relies heavily on a semantic understanding of the environment through visual feedback [29]. Despite its importance, embedding information such as traversability through this modality has its unique challenges [52]. First, visual observations (e.g., RGB images) from robots are represented in high-dimensional tensors, requiring exponentially more training data. Also, small changes in environments coming from illumination, weather, or time of day may result in catastrophic failure of machine learning models [12].
Task. We construct our task scene using black and white colored foam tiles (see Fig. 7). These tiles are inexpensive, safe, and widely available. Surrounded by black tiles, a “figure-8” path is laid using white tiles. The platform must remain on the white path while avoiding the surrounding black tiles. A dynamic version of this task is also used in evaluation; black “barricade” tiles are dynamically removed and placed to demonstrate robustness and real-time navigation reasoning.
We generate simulation environments with black and white areas as shown in Fig. 8 to train the visual policy. Each example in the figure represents each sub-environment, and our entire map is composed of sub-environments. This allows us to assign different difficulties for each sub-environment. More details of map generation can be found in Appendix B-C.
Evaluation. We measure success or failure over the number of five trials. We define success if the robot finishes at least one lap without staying in one place for more than 5 seconds.
@@
Results. For quantitative evaluation, we run experiments five times and count the number of successes for each setting. As reported in Table III, image augmentations play an essential role in Sim2Real generalization. The policies without augmentation fail at generalization. Also, despite the better learning capability of CNNs, they result in worse generalization to real. We found CNN training often collapses to a trivial solution with the reward designs remaining the same as MLP trainings. Such results imply that CNNs may overfit to simulated data and require more careful design in rewards and loss functions for stable training.
We qualitatively evaluate the results through recorded videos in the supplementary materials. While both policies with augmentation predict meaningful actions, the -based policy generates more smooth and precise actions. On the other hand, the CNN-based policy often generates noisy actions as shown in fast steering changes and motor cogging noise.
VI Conclusion
In modern robotics, state-of-the-art methods are focusing more on complex, costly, and proprietary platforms and ecosystems, leaving behind more resource-constrained audiences. Wheeled platforms, on the other hand, already have deep roots in the broader community of educators, students, and enthusiasts, often acting as the entry-level platform of choice for beginning roboticists.
@@ VI Conclusion
To kickstart this integration, we demonstrate the entire Sim2Real pipeline on three state-of-the-art policy types. Each policy and task is motivated by their fundamental relevance to general robot learning and Sim2Real. Drifting is a complex control problem highlighting accurate, randomized, and massive physical simulation. Elevation traversal defines safe behavior at the interface between perception and control, highlighting physical scene generation and massive interaction. Visual navigation integrates a sensor modality with substantial research attention and pre-training potential while highlighting challenges in data diversity and transferability.
Wheeled Lab is entirely low-cost and open-source. Experiment videos, code, and more information can be found on our website.
VII Limitations
While an objective of Wheeled Lab is to support education, this work has not yet run a study to evaluate how comprehensible the pipeline is to robotics students. Similarly to F1Tenth, we aim to construct an open curriculum around Wheeled Lab to help introduce difficult or interdisciplinary topics that may be sources of confusion. This study will help reveal what steps in the Sim2Real process might be most challenging or unfamiliar.
While iterative testing for unique conditions can serve pedagogical purposes, the appeal of building, training, and deploying a platform that “just works” can be inspirational for beginner roboticists. However, the policies presented here in their current open-source state must be met with patience. Thus, custom low-cost wheeled platforms invite yet another frontier of modern robot learning: adaptation [7]. For instance, the development of a drifting pipeline which adequately adapts to friction on deployment can help lower the barrier-to-entry even further.
Due to the lack of access to facilities with meaningful elevation features, our evaluation of $\boldsymbol{\pi}_{\text{elev}}$ used a motion capture system, a comparatively inaccessible piece of equipment for our target audience. We aim to establish a benchmark in the future which still evaluates elevation-based training using onboard compute and sensing while remaining accessible to individuals in a variety of geographical regions.
Acknowledgments
TH is supported by the NSF GRFP under Grant No. DGE 2140004.
@@ Appendix B Design and Implementation Details
We recommend interested readers view the configuration files for each task found in GitHub (https://github.com/UWRobotLearning/WheeledLab) for a detailed overview. Isaac Lab environment configurations primarily include parameter settings and reward functions and can be quickly parsed while efficiently communicating environment behavior. We highlight some notable details in this section.
B-A Drifting
@@ B-A3 Articulations
Suspension helps close the gap on load shifting during turns. With a suspension travel of 4 cm, these shifts have consequences in the total friction and “thrust” produced while throttling through turns. Interestingly, suspension articulation also helped to stabilize the simulation. At a wheel speed of 3 m/s, the rotation rate of the wheels reaches 60 rad/s, occasionally resulting in collision penetrations during simulation which cause unrealistic “jumping” through drifted turns. Damped suspension joints helped to stabilize these collisions without requiring smaller simulation time steps which would significantly affect training times.
B-A4 Rewards
@@
    Cross-Track Distance. The lateral distance of the vehicle from an oval track line. This is defined without time-dependent reference trajectories as often done in model-based approaches or even in deep RL methods [9]. This is highly weighted otherwise the agent tends to take larger, safer turns.
@@
    Side-Slip. Only stable side-slip angles rewarded. Stable side-slip angles are manageable through the steering limits (0-30).
@@
    Turn-Energy. Shaping term for increasing velocity rewards specifically when inside turn regions.
@@
    Turn-Left-Go-Right. Shaping term for encouraging exploration to discover an alternate turning mode (e.g. drifting) by counter-steering. Strictly positive when angular velocity and steering command are opposite to each other. 0 otherwise.
@@ B-B1 Training
B-B2 Suspension
As mentioned in Appendix B-A, suspension joints were added to the open-source MuSHR articulations. This is critical for elevation where maintaining momentum up a ramp helps training stability. Without suspension articulation, climbing ramps have an abrupt and destabilizing impact at higher speeds.
B-B3 Rewards
@@
    Goal Velocity. Rewards velocity projected onto goal heading vector.
@@ B-C Visual
B-C1 Visual map generation
We designed the problem by encoding traversability into color information. Black tiles are not-traversable regions while white indicates safe to travel. We divided the whole map into multiple sub-environments that are individually configurable. Then, we randomized white paths over the black area in each sub-environment. Specifically, each sub-environment comprises 9 cells, and we sample a start point in each cell and generate paths toward random endpoints in the sub-environment. We made the number of random paths from a single start point configurable, allowing the users to adjust the difficulty of the training (see Fig. 8). This can later be used for curriculum learning by assigning different difficulties for sub-environments. We tested the generated map with more than 2,000 agents in parallel running seamlessly with a single GPU. The algorithm overview for sub-environment generation is presented in Alg. 1.
@@ B-C1 Visual map generation
2:$(H_{\text{env}},W_{\text{env}})\leftarrow$ of rows and # of columns for each sub environment
3:$(H_{\text{cell}},W_{\text{cell}})\leftarrow$ of rows and # of columns for each cell
4:$N_{\text{walkers}}\leftarrow$ of walkers for path generation
@@
7$\triangleright$ Initialize traversability map; 0 - not traversable / 1 - traversable
8:$\mathbf{T}\in\{0,1\}^{H_{\text{env}}\times W_{\text{env}}}\leftarrow\mathrm{\mathbf{0}}$
@@
10$\triangleright$ Sample start points from each cell
11:$\mathbf{P}^{\text{start}}\leftarrow[\,]$triangleright Initialize an array to save start points
12$\triangleright$ Sample random start points for each cell
@@
17:        $\mathbf{P}^{\text{start}}\,\mathrel{+}=[(p_{\text{row}},p_{\text{col}})]$triangleright Append the start point to $\mathbf{P}^{\text{start}}$
@@
29:        GenerateRandomPaths$(\mathbf{T},\,(p_{\text{row}},p_{\text{col}}),\,(q_{\text{row}},q_{\text{col}}))$
@@
B-C2 Model Architectures
We adopt multi-layer perceptron (MLP) as our basic policy network. For ablations, we add a simple 3-layer convolutional neural network (CNN) that serves as an image feature extractor before the policy networks. The extracted features are concatenated to the rest of the observation terms such as linear/angular velocities and last action and fed into the MLP policy networks. We observe that inducing CNN takes longer training time to learn optimal actions and often collapses into a trivial solution - agents decide to spin in a circle, only maximizing the velocity rewards.
B-C3 Image Augmentations
@@ B-C4 Rewards
B-C5 Terminal Conditions
We experimented with different terminations. In addition to the standard time-out and out-of-bounds terminal conditions, we added a terminal condition when the robot is on black. However, this often resulted in a lack of exploration, and policies that produced “jerky” and undesirable actions, such as idling at the initial location.

### v2

@@ 4.2 Elevation
Background. On unstructured and uneven terrain, an elevation map is essential for identifying traversable areas in the scene. However, traversability is strongly dependent on agent morphology and dynamic state. Along with the ground geometry, the following affect how to safely traverse uneven terrain: the platform’s ground clearance, center-of-mass height, wheel size, maximum torque, suspension, and momentum. Poor traversal can result in catastrophic failures, such as rollovers, and motor (or engine) stalling.
@@
Task. To show that the trained elevation-based policy is capable of geometric reasoning, we construct a scene with two types of elevation features that contribute to the overall elevation map, i.e., walls and ramps. Crucially, the incorporation of ramps sets the task apart from sole obstacle avoidance since ramps are features with height but are still traversable with the correct approach strategy. The starting and goal positions are placed on opposite sides of the testing area, with the elevation features placed in between. The platform must safely traverse the scene to arrive at the goal position by avoiding obstacles and ascending ramps when locally feasible.
@@
Setup. States provided to the policy are goal-relative position, orientation, and a local ($2.5\times 2.5$ meters) body-centric elevation map. A local map is sampled from a prior global elevation map. Although the platform is capable of creating onboard maps [1], a configurable indoor environment requires higher mapping resolution than is available for this evaluation.
@@
Results. We find that the baseline primarily deviates from any elevation features to approach the goal. When the baseline does ascend the ramp, it quickly falls off the sides. However, $\boldsymbol{\pi}_{\text{elev}}$ can both traverse the ramp safely and navigate through subsequent obstacles to reach the goal. Like $\overline{\boldsymbol{\pi}}_{\text{drift}}$, the lower number of simulated agents gives the baseline policy significantly fewer opportunities to explore correct trajectories in a narrow range of safe states.
4.3 Visual
@@
Figure 6: Captured trajectories of visual policy in real experiments. The trajectories are obtained from the motion capture system. We include videos of each setting and trial in the supplementary.
Background. Visual navigation traditionally relies heavily on a semantic understanding of the environment through visual feedback [51]. Despite its importance, embedding information such as traversability through this modality has its unique challenges [52]. First, visual observations (e.g., RGB images) from robots are represented in high-dimensional tensors, requiring exponentially more training data. Also, small changes in environments that result from illumination, weather, or time of day can result in catastrophic failure of machine learning models [53].
Task. We construct our task scene using black and white foam tiles (see Fig. 6). These tiles are inexpensive, safe, and widely available. Surrounded by black tiles, a “figure-8” path is laid using white tiles. The platform must remain on the white path while avoiding the surrounding black tiles. A dynamic version of this task is also used in evaluation; black “barricade” tiles are dynamically removed and placed to demonstrate robustness and real-time navigation reasoning.
We generate simulation environments with black and white areas, as we show in Fig. 7 to train the visual policy. Each example in the figure represents a sub-environment, and our entire map is composed of sub-environments. This allows us to assign different difficulties for each sub-environment. See Appendix B.3 for additional details about map generation.
Evaluation. We measure success or failure over five trials. We define success as the robot finishing at least one lap without staying in one place for more than 5 seconds.
@@
Table 3: Visual policy results. Results demonstrate the effectiveness of image augmentations for Sim2Real generalization. While we expect CNN to be more capable of learning, MLP presents better generalization capability. We include videos of each run in the supplementary materials.
@@
Results. For quantitative evaluation, we run experiments five times and count the number of successes for each setting. As reported in Table 3, image augmentation is essential in Sim2Real generalization. The policies without augmentation fail at generalization. Further, despite the better learning capability of CNNs, they result in worse generalization to the real world; we also found that CNN training often collapses to a trivial solution, with the reward designs remaining the same as MLP training. Such results imply that CNNs may overfit to simulated data and require more careful design in rewards and loss functions for stable training.
In the supplementary materials, we qualitatively evaluate the results using recorded videos. Although both policies with augmentation predict meaningful actions, the MLP-based policy generates smoother and precise actions. In contrast, the CNN-based policy often generates noisy actions, as shown in fast steering changes and audible motor cogging.
5 Conclusion
Drawing from our results in Section 4, observe that there is an alarming gap in techniques between the scientific and broader communities. Methods and modalities (e.g., parallelization, corruption, elevation), almost native in popular implementations and frameworks for platforms like humanoids and quadrupeds, are scattered throughout different implementations and research efforts for wheeled platforms.
@@ 5 Conclusion
To kickstart this integration, we demonstrate the entire Sim2Real pipeline on three state-of-the-art policy types. Each policy and task is motivated by their fundamental relevance to general robot learning and Sim2Real. Drifting is a complex control problem that highlights accurate, randomized, and massive physical simulation. Through Wheeled Lab, we demonstrate the first zero-shot transfer of a drifting policy thus far in the literature. Elevation traversal defines safe behavior at the interface between perception and control, highlighting physical scene generation and massive interaction. Visual navigation integrates a sensor modality with substantial research attention and pre-training potential while showcasing challenges in data diversity and transferability.
Wheeled Lab is entirely low-cost and open-source. Please refer to the project website (https://uwrobotlearning.github.io/WheeledLab/) for experiment videos and more information.
6 Limitations
Although an objective of Wheeled Lab is to support education, this work has not yet conducted a study to evaluate how comprehensible the pipeline is to robotics students. Like F1Tenth, we aim to construct an open curriculum around Wheeled Lab to help introduce difficult or interdisciplinary topics that may be sources of confusion. This study will help reveal what steps in the Sim2Real process might prove most challenging or unfamiliar.
While iterative testing for unique conditions can serve pedagogical purposes, the appeal of building, training, and deploying a platform that “just works” can be inspirational for beginner roboticists. However, the policies presented here in their current open-source state must be met with patience and an intent to iterate. Thus, custom low-cost wheeled platforms invite yet another frontier of modern robot learning: robust adaptation [54]. For instance, the development of a drifting pipeline that adequately adapts to friction on deployment can help lower the barrier-to-entry even further.
Towards Accessible Sim2Real
The development of Wheeled Lab is constrained by both platform and development to low-cost yet easily obtainable resources. Mechanical hysteresis and tolerances of hobby-grade components, limited onboard compute and sensing, visual diversity, methodological simplicity, access to evaluation terrain, access to open-source packages, and access to training resources are all considered in the development of Wheeled Lab. Sim2Real accessibility is not only a challenging constraint on robot hardware but on the entire development process.
In accessibility efforts, experts are perhaps paradoxically limited by their expertise, making viewing problems from the perspective of a novice or non-traditional practitioner challenging or impossible. We find that broader audiences tend to benefit significantly from approachable infrastructure, familiar abstractions, modularity, and straightforward, reproducible, yet instructive tasks and experiments. Wheeled Lab has fortunately garnered interest from the broader community in support of these objectives, receiving contributions and feedback from the perspective of non-scientific audiences. Meanwhile, we expect that experts are better positioned to contribute fundamental concepts with vision. Wheeled Lab is an essential step towards aligning these communities and fostering a future of robot learning where all roboticists can equip themselves with indispensable skills in hardware, deployment, and iterative development.
@@ Appendix B Design and Implementation Details
Isaac Lab environment configurations primarily include parameter settings and reward functions, and can be quickly parsed while efficiently communicating environment behavior. We highlight some notable details in this section.
B.1 Drifting
@@ B.1.3 Articulations
Suspension helps close the gap on load shifting during turns. With a suspension travel of 4 cm, these shifts have consequences in the total friction and “thrust” produced while throttling through turns. Interestingly, suspension articulation also helped to stabilize the simulation. At a wheel speed of 3 m/s, the rotation rate of the wheels reaches 60 rad/s, occasionally resulting in collision penetrations during simulation which cause unrealistic “jumping” through drifted turns. Damped suspension joints helped to stabilize these collisions without requiring smaller simulation time steps which would significantly lengthen training times.
B.1.4 Rewards
@@
    Cross-Track Distance. The lateral distance of the vehicle from an oval track line. This is defined without time-dependent reference trajectories as often done in model-based approaches or even in deep RL methods [30]. This is highly weighted otherwise, the agent tends to take larger, safer turns.
@@
    Side-Slip. Only stable side-slip angles were rewarded. Stable side-slip angles are manageable through the steering limits (0-30).
@@
    Turn-Energy. Shaping term for increasing velocity rewards, specifically when inside turn regions.
@@
    Turn-Left-Go-Right. Shaping term for encouraging exploration to discover an alternate turning mode (e.g., drifting) by counter-steering. Strictly positive when the angular velocity and the steering command are opposite to each other. 0 otherwise.
@@ B.2.1 Training
B.2.2 Suspension
As mentioned in Appendix B.1, suspension joints were added to the open-source MuSHR articulations. This is critical for elevation where maintaining momentum up a ramp helps training stability. Without suspension articulation, climbing ramps have an abrupt and destabilizing impact at higher speeds.
B.2.3 Rewards
@@
    Goal Velocity. Rewards velocity projected onto the goal heading vector.
@@ B.3 Visual
B.3.1 Visual map generation
We designed the problem by encoding traversability into color information. Black tiles are non-traversable regions while white indicates safe to travel. We divided the whole map into multiple sub-environments that are individually configurable. Then, we randomized white paths over the black area in each sub-environment. Specifically, each sub-environment comprises 9 cells, and we sample a start point in each cell and generate paths toward random endpoints in the sub-environment. We made the number of random paths from a single start point configurable, allowing the users to adjust the difficulty of the training (see Fig. 7). This can later be used for curriculum learning by assigning different difficulties for sub-environments. We tested the generated map with more than 2,000 agents in parallel, running seamlessly with a single Nvidia RTX 3080 GPU. The algorithm overview for sub-environment generation is presented in Alg. 1.
@@ B.3.1 Visual map generation
[…]
@@
[…]
@@
[…]
@@
17:   $\mathbf{P}^{\text{start}}\,\mathrel{+}=[(p_{\text{row}},p_{\text{col}})]$ $\triangleright$ Append the start point to $\mathbf{P}^{\text{start}}$
@@
[…]
@@
B.3.2 Model Architectures
We adopt a multi-layer perceptron (MLP) as our basic policy network. For ablations, we add a simple 3-layer convolutional neural network (CNN) that serves as an image feature extractor before the policy networks. The extracted features are concatenated to the rest of the observation terms, such as linear/angular velocities and last action, and fed into the MLP policy networks. We observe that inducing CNN takes longer training time to learn optimal actions and often collapses into a trivial solution - agents decide to spin in a circle, only maximizing the velocity rewards.
B.3.3 Image Augmentations
@@ B.3.4 Rewards
B.3.5 Terminal Conditions
We experimented with different terminations. In addition to the standard time-out and out-of-bounds terminal conditions, we added a terminal condition when the robot is on black. However, this often resulted in a lack of exploration, and policies that produced “jerky” or undesirable actions, such as idling at the initial location.